# On the SOS Rank of Simple and Diagonal Biquadratic Forms

### v1

@@ Abstract
We study the sum-of-squares (SOS) rank of simple and diagonal biquadratic forms. For simple biquadratic forms in $3\times 3$ variables, we show that the maximum SOS rank is exactly $6$, attained by a specific six-term form. We further prove that for any $m\geq 3$, there exists an $m\times m$ simple biquadratic form whose SOS rank is exactly $2m$, providing a general lower bound that extends the $3\times 3$ case. For diagonal biquadratic forms with nonnegative coefficients, we prove an SOS rank upper bound of $7$, improving the general bound of $8$ for $3\times 3$ forms. In addition, we extend the techniques to a broader class of sparse biquadratic forms, obtaining combinatorial upper bounds and constructing explicit families whose SOS rank grows linearly with the number of terms. These results provide new lower and upper bounds on the worst-case SOS rank of biquadratic forms and highlight the role of structure in reducing the required number of squares.
Keywords. Biquadratic forms, sum-of-squares, SOS rank, simple forms, diagonal forms, sparse forms, positive semidefinite.
@@ 1 Introduction
These results provide concrete bounds for two structured families and highlight the interplay between combinatorial pattern (simple forms) and coefficient structure (diagonal forms).
These results provide concrete bounds for two structured families and highlight the interplay between combinatorial pattern (simple forms) and coefficient structure (diagonal forms). In Section 4 we extend the methods to sparse biquadratic forms, obtaining combinatorial upper bounds and exhibiting families whose SOS rank equals the number of terms, showing that the techniques developed here apply to a wider class of structured forms.
@@ 3 Diagonal Biquadratic Forms
4 Sparse Biquadratic Forms
The methods developed for simple and diagonal forms extend naturally to biquadratic forms whose nonzero coefficients are confined to a small subset of all possible monomials. We call such forms sparse and show that their SOS rank can be bounded from above by combinatorial properties of the support, and from below by algebraic orthogonality arguments analogous to those in Section 2.
4.1 Definitions and notation
Let
$P({\bf x},{\bf y})=\sum_{i,k=1}^{m}\sum_{j,l=1}^{n}a_{ijkl}\,x_{i}x_{k}y_{j}y_{l}$
be a biquadratic form. Define its $x$-support and $y$-support as
$\mathcal{S}_{x}(P)=\{(i,k):a_{ijkl}\neq 0\text{ for some }j,l\},\qquad\mathcal{S}_{y}(P)=\{(j,l):a_{ijkl}\neq 0\text{ for some }i,k\}.$
We say $P$ is $(\alpha,\beta)$-sparse if $|\mathcal{S}_{x}(P)|\leq\alpha$ and $|\mathcal{S}_{y}(P)|\leq\beta$.
Example. A simple biquadratic form (only terms $x_{i}^{2}y_{j}^{2}$) has $\mathcal{S}_{x}=\{(i,i):i=1,\dots,m\}$, $\mathcal{S}_{y}=\{(j,j):j=1,\dots,n\}$; hence it is $(m,n)$-sparse.
4.2 A combinatorial extraction lemma
The key tool for upper bounds is the detection of $T$-subforms inside a sparse form.
Lemma 4.1 (Rectangle completion).
Suppose $P$ is a PSD biquadratic form whose support contains indices $(i,i),(i,i^{\prime}),(i^{\prime},i),(i^{\prime},i^{\prime})$ in $\mathcal{S}_{x}$ and $(j,j),(j,j^{\prime}),(j^{\prime},j),(j^{\prime},j^{\prime})$ in $\mathcal{S}_{y}$ with $i\neq i^{\prime},\;j\neq j^{\prime}$. Then $P$ contains a positive?definite subform (after scaling) isomorphic to
$T_{iji^{\prime}j^{\prime}}=x_{i}^{2}y_{j}^{2}+x_{i^{\prime}}^{2}y_{j^{\prime}}^{2}+x_{i}^{2}y_{j^{\prime}}^{2}+x_{i^{\prime}}^{2}y_{j}^{2},$
which satisfies $\operatorname{sos}(T_{iji^{\prime}j^{\prime}})=2$.
@@ 4.2 A combinatorial extraction lemma
Because $P$ is PSD, the coefficients $a_{iijj},a_{i^{\prime}ij^{\prime}j^{\prime}},a_{iij^{\prime}j^{\prime}},a_{i^{\prime}ijj^{\prime}}$ are nonnegative. Since all four index pairs belong to the support, these coefficients are positive. Define
$\tilde{T}=\frac{a_{iijj}}{c}\,x_{i}^{2}y_{j}^{2}+\frac{a_{i^{\prime}ij^{\prime}j^{\prime}}}{c}\,x_{i^{\prime}}^{2}y_{j^{\prime}}^{2}+\frac{a_{iij^{\prime}j^{\prime}}}{c}\,x_{i}^{2}y_{j^{\prime}}^{2}+\frac{a_{i^{\prime}ijj^{\prime}}}{c}\,x_{i^{\prime}}^{2}y_{j}^{2},$
where $c>0$ is a common scaling factor. Then $\tilde{T}$ is a positive multiple of $T_{iji^{\prime}j^{\prime}}$, and by (2.1) it can be written as two squares. Hence any SOS decomposition of $P$ must use at least two squares for this subform. ∎
4.3 Upper bounds from support size
Theorem 4.2 (Sparsity upper bound).
Let $P$ be a PSD biquadratic form that is $(\alpha,\beta)$-sparse. Then
$\operatorname{sos}(P)\leq\min\!\bigl\{\alpha\beta,\;\bigl\lceil\tfrac{\alpha\beta}{2}\bigr\rceil\bigr\}.$
Proof.
The first bound $\alpha\beta$ is trivial: each of the at most $\alpha\beta$ possible term positions could be written as a separate square.
For the second bound, observe that if the support contains a $2\times 2$ rectangle as in Lemma ??, we can cover four coefficient positions with two squares instead of four. In the worst case (maximal packing), we can pair up to $\lfloor\alpha\beta/2\rfloor$ disjoint such rectangles, each needing two squares, plus at most one leftover single term needing one square. Thus the total number of squares is at most $\lceil\alpha\beta/2\rceil$. ∎
Remark. The bound $\lceil\alpha\beta/2\rceil$ is often pessimistic because disjoint rectangles may not exist; the true maximum lies between the two given bounds.
4.4 A sparse family with high SOS rank
The next construction shows that sparsity alone does not drastically reduce the SOS rank — it can still be proportional to the number of terms.
Theorem 4.3 (Sparse high?rank family).
For $m\geq 3$ define the double?chain form
$D_{m}({\bf x},{\bf y})=\sum_{i=1}^{m}x_{i}^{2}y_{i}^{2}+\sum_{i=1}^{m-1}x_{i}^{2}y_{i+1}^{2}+\sum_{i=1}^{m-1}x_{i+1}^{2}y_{i}^{2}.$
Then (i) $D_{m}$ is (2m1, 2m1)(2m-1,\,2m-1)-sparse and has 3m23m-2 terms; (ii) $\operatorname{sos}(D_{m})=3m-2$.
Proof.
(i) The $x$-support is $\{(i,i):i=1,\dots,m\}\cup\{(i,i+1),(i+1,i):i=1,\dots,m-1\}$, which has size m+2(m1)=2m1m+2(m-1)=2m-1; similarly for $\mathcal{S}_{y}$. There are exactly 3m23m-2 square terms.
(ii) Upper bound: Write each term as a separate square, giving $\operatorname{sos}(D_{m})\leq 3m-2$.
Lower bound: Suppose $D_{m}=\sum_{t=1}^{R}L_{t}^{2}$ with $L_{t}=\sum_{i,j}c_{ij}^{(t)}x_{i}y_{j}$. Because $D_{m}$ contains only square terms $x_{a}^{2}y_{b}^{2}$, for each $(a,b)$ in the support $S$ we have
$\sum_{t=1}^{R}\bigl(c_{ab}^{(t)}\bigr)^{2}=1,$
and $c_{ij}^{(t)}=0$ for $(i,j)\notin S$. For cross terms $x_{i}x_{p}y_{j}y_{q}$ with $(i,j)\neq(p,q)$ to vanish, we obtain, exactly as in the proof of Theorem 2.7, that the vectors $\mathbf{C}_{ab}=(c_{ab}^{(1)},\dots,c_{ab}^{(R)})$ for $(a,b)\in S$ are pairwise orthogonal in $\mathbb{R}^{R}$. Since |S|=3m2|S|=3m-2, orthogonality forces $R\geq 3m-2$. ∎
Thus even with relatively small supports $(\alpha,\beta)\approx 2m$, the SOS rank can be as large as the number of terms, showing that the trivial upper bound $\alpha\beta$ is sometimes tight.
4.5 Link to the $3\times 3$ simple forms
The $3\times 3$ simple forms studied in Section 2 are a special case of sparse forms with $\mathcal{S}_{x}=\{(1,1),(2,2),(3,3)\}$ and $\mathcal{S}_{y}=\{(1,1),(2,2),(3,3)\}$, i.e., they are $(3,3)$-sparse. Theorems 2.1–2.5 can be viewed as a complete solution for this particular support pattern.
Corollary 4.4 (Sparsity interpretation of the $3\times 3$ case).
Let $P$ be a $3\times 3$ simple biquadratic form (hence $(3,3)$-sparse). Then $\operatorname{sos}(P)\leq 6$, and the bound is attained by $P_{3,3,6}$. Moreover, if $P$ contains at least 7 terms, Lemma ?? guarantees the presence of a $T$-subform, which immediately yields the improved bounds of Proposition 2.3.
Proof.
For a $3\times 3$ simple form, $\alpha=\beta=3$, so Theorem 4.2 gives $\operatorname{sos}(P)\leq\lceil 9/2\rceil=5$ — but this is weaker than the true maximum 6 because the support does not contain enough disjoint rectangles. The exact maximum 6 follows from the specific combinatorics of the $3\times 3$ grid, as proved in Section 2. The appearance of a $T$-subform when at least 7 terms are present is exactly Lemma 2.2, which is a special case of Lemma ??. ∎
This corollary shows that the results of Section 2 fit naturally into the broader sparse?form framework, and that the combinatorial rectangle argument is the common engine behind both the $3\times 3$ simple?form bounds and the more general sparse bounds.
4.6 Connection to Gram matrix sparsity
A biquadratic form $P$ is SOS iff its Gram matrix $G\in\mathbb{R}^{mn\times mn}$ (with rows indexed by $(i,j)$ and columns by $(k,l)$, entries $a_{ijkl}$) is PSD and has rank $R=\operatorname{sos}(P)$. Sparsity of $P$ induces a block?sparse pattern in $G$: the submatrix with rows $(i,j),(i^{\prime},j^{\prime})$ and columns $(k,l),(k^{\prime},l^{\prime})$ is nonzero only if $(i,k),(i^{\prime},k^{\prime})\in\mathcal{S}_{x}$ and $(j,l),(j^{\prime},l^{\prime})\in\mathcal{S}_{y}$. Lemma ?? corresponds to the existence of a fully dense $2\times 2$ principal submatrix of $G$, forcing that submatrix to have rank at least 2. Thus, lower bounds on $\operatorname{sos}(P)$ can be derived from clique numbers in the graph whose edges correspond to such fully dense $2\times 2$ blocks — a purely combinatorial invariant of the support.
4.7 Open problems
  1. 1.
    Tight sparse bound. For fixed $\alpha,\beta$, determine
    $\max\{\operatorname{sos}(P):P\text{ is PSD and }(\alpha,\beta)\text{-sparse}\}.$
    Theorem ?? shows it can reach $\alpha+\beta-1$; is that the maximum?
  2. 2.
    Random sparse supports. If each pair $(i,k)$ is included in $\mathcal{S}_{x}$ independently with probability $p_{x}$, and similarly for $\mathcal{S}_{y}$, what is the expected SOS rank of a random PSD form with that support? The appearance of $T$-subforms becomes a problem of detecting rectangles in random bipartite graphs.
  3. 3.
    Algorithmic exploitation. Can one use the combinatorial detection of $T$-subforms to reduce the size of semidefinite programs when solving sparse SOS problems? A preprocessing step that replaces each detected $T$-block with two squares could lower the computational cost.
These results illustrate that the techniques developed for simple and diagonal forms — combinatorial rectangle arguments, orthogonal?vector methods, and explicit high?rank constructions — extend effectively to sparse biquadratic forms, offering new bounds and highlighting the interplay between support structure and SOS complexity.
@@ 5 Concluding Remarks
We have determined the exact maximum SOS rank ($6$) for $3\times 3$ simple biquadratic forms and proved an improved upper bound ($7$) for $3\times 3$ diagonal forms. Moreover, we have shown that for every $m\geq 3$ there exists an $m\times m$ simple biquadratic form whose SOS rank equals $2m$, extending the $3\times 3$ case and providing a linear lower bound in $m$. The extension to sparse biquadratic forms (Section 4) demonstrates that the methods developed here apply to a wider class of structured forms, yielding combinatorial upper bounds and families whose SOS rank grows with the number of terms.
A natural open problem is to determine the exact value of $\mathrm{BSR}(3,3)$, the maximum SOS rank among all $3\times 3$ SOS biquadratic forms. Our results give $6\leq\mathrm{BSR}(3,3)\leq 8$; closing this gap would be an interesting next step. Another direction is to study whether the $2m$ lower bound for simple $m\times m$ forms is optimal, or whether even higher SOS ranks can be achieved with different supports $S$. The sparse form framework introduced in Section 4 offers a systematic way to attack such questions by linking SOS rank to the combinatorial geometry of the support.

### v2

@@ Abstract
We study the sum-of-squares (SOS) rank of simple and diagonal biquadratic forms. For simple biquadratic forms in $3\times 3$ variables, we show that the maximum SOS rank is exactly $6$, attained by a specific six-term form. We further prove that for any $m\geq 3$, there exists an $m\times m$ simple biquadratic form whose SOS rank is exactly $2m$, providing a general lower bound that extends the $3\times 3$ case. For diagonal biquadratic forms with nonnegative coefficients, we prove an SOS rank upper bound of $7$, improving the general bound of $8$ for $3\times 3$ forms. These results provide new lower and upper bounds on the worst-case SOS rank of biquadratic forms and highlight the role of structure in reducing the required number of squares.
Keywords. Biquadratic forms, sum-of-squares, SOS rank, simple forms, diagonal forms, positive semidefinite.
@@ 1 Introduction
Beyond structured forms, our work also provides evidence for the exact value of the maximum SOS rank among all $3\times 3$ biquadratic forms. We conjecture that $\mathrm{BSR}(3,3)=6$ (Conjecture 4.2), a value attained by the simple form $P_{3,3,6}$; we support this by showing that enriching this form with certain additional squares does not increase its SOS rank.
@@ 3 Diagonal Biquadratic Forms
4 Toward the Maximum SOS Rank for $3\times 3$ Forms
The simple form $P_{3,3,6}$ has SOS rank $6$, providing the lower bound $\mathrm{BSR}(3,3)\geq 6$. To investigate whether $6$ might also be an upper bound, we examine what happens when we enrich $P_{3,3,6}$ with an additional square term.
A natural candidate for a form that could potentially increase the SOS rank is
$P_{+}({\bf x},{\bf y})=P_{3,3,6}({\bf x},{\bf y})+(x_{1}y_{3}+x_{2}y_{1})^{2}.$
Notice that $P_{3,3,6}$ contains neither $x_{1}^{2}y_{3}^{2}$ nor $x_{2}^{2}y_{1}^{2}$. Thus $P_{+}$ introduces two new square terms ($x_{1}^{2}y_{3}^{2}$ and $x_{2}^{2}y_{1}^{2}$) together with the cross term $2x_{1}x_{2}y_{1}y_{3}$. This enrichment fills missing entries in the support of $P_{3,3,6}$ and creates new bilinear couplings; therefore, if any added square were to force an increase in the SOS rank, this would be a likely example.
A priori we have $\operatorname{sos}(P_{+})\leq 7$, because we may simply take the six separate squares of $P_{3,3,6}$ together with the explicit square $(x_{1}y_{3}+x_{2}y_{1})^{2}$. Surprisingly, a more efficient decomposition exists.
Proposition 4.1.
$\operatorname{sos}(P_{+})\leq 6$.
@@ 4 Toward the Maximum SOS Rank for $3\times 3$ Forms
The following six squares give an explicit SOS decomposition:
$P_{+}({\bf x},{\bf y})$ $=\Bigl(\tfrac{1}{2}x_{1}y_{1}+x_{2}y_{3}+\tfrac{\sqrt{3}}{2}x_{1}y_{2}\Bigr)^{2}$
$\quad+\Bigl(\tfrac{1}{2}x_{2}y_{1}+x_{1}y_{3}-\tfrac{\sqrt{3}}{2}x_{2}y_{2}\Bigr)^{2}$
$\quad+\Bigl(\tfrac{\sqrt{3}}{2}x_{1}y_{1}-\tfrac{1}{2}x_{1}y_{2}\Bigr)^{2}$
$\quad+\Bigl(\tfrac{1}{2}x_{2}y_{2}+\tfrac{\sqrt{3}}{2}x_{2}y_{1}\Bigr)^{2}$
$\quad+(x_{3}y_{1})^{2}+(x_{3}y_{3})^{2}.$
Expanding and simplifying verifies that all terms match $P_{3,3,6}+(x_{1}y_{3}+x_{2}y_{1})^{2}$, with all unwanted cross terms canceling. ∎
Thus, even in this “dangerous” case, where two missing square terms are added together with a new cross term - the SOS rank does not exceed $6$. This observation lends further support to the conjecture that no $3\times 3$ biquadratic form requires more than six squares.
Conjecture 4.2.
For $3\times 3$ biquadratic forms,
$\mathrm{BSR}(3,3)=6.$
Remark 4.3.
Proposition 4.1 shows even in the “dangerous” case, the SOS rank does not exceed $6$. Together with Theorem 2.5, which establishes 66 as the maximum SOS rank for simple forms, this strongly supports Conjecture 4.2. The currently best known bounds are $6\leq\mathrm{BSR}(3,3)\leq 8$; closing this gap remains an intriguing open problem.
@@ 5 Concluding Remarks
We have determined the exact maximum SOS rank ($6$) for $3\times 3$ simple biquadratic forms and proved an improved upper bound ($7$) for $3\times 3$ diagonal forms. Moreover, we have shown that for every $m\geq 3$ there exists an $m\times m$ simple biquadratic form whose SOS rank equals $2m$, extending the $3\times 3$ case and providing a linear lower bound in $m$. These results demonstrate how structural constraints can either lower or raise the required number of squares in an SOS decomposition.
A natural open problem is to determine the exact value of $\mathrm{BSR}(3,3)$, the maximum SOS rank among all $3\times 3$ SOS biquadratic forms. Our results give $6\leq\mathrm{BSR}(3,3)\leq 8$, and we conjecture that in fact $\mathrm{BSR}(3,3)=6$ (Conjecture 4.2). Proving or disproving this would complete the picture for $3\times 3$ biquadratic forms and remains an interesting next step.